# A Mathematical Framework for Misinformation Propagation in Complex Networks: Topology-Dependent Distortion and Control

### v2

@@ VI.3 Barabási–Albert Networks
Scale-free networks generated by the Barabási–Albert (BA) preferential attachment model barabasi_science_1999 provide a natural framework for examining the role of heterogeneous connectivity in misinformation. In this model, each newly added node attaches to $m$ existing nodes with probability proportional to their degrees, producing a power-law degree distribution
@@
The biological relevance of scale-free architecture is well documented. Metabolic networks jeong2000; wagner2001small, gene-regulatory circuits lee2002transcriptional; babu2004structure, and protein–protein interaction networks jeong_nature_2001; han2004evidence all display degree exponents between 2 and 3. In such systems, hubs frequently correspond to essential components—such as ATP in metabolism or master transcriptional regulators—that coordinate activity across multiple pathways and support robust communication. For context, Table 2 compares several biological scale-free networks with a BA network of matched average degree. Although empirical exponents vary across the range 2–3, the BA model consistently produces an exponent close to 3.
@@ VII.2 Translation of the perceived distribution
To understand why Eq. (31) preserves the functional shape of the distribution, consider the case where the true state $X_{i}$ is Gaussian. The perceived value $X^{(j\leftarrow i)}$ is the sum of this Gaussian variable, a constant drift, and a centered binomial fluctuation. Because the fluctuation term is a sum of independent Bernoulli increments, it is approximately Gaussian by the central limit theorem feller_prob; and the sum of two Gaussian variables is Gaussian papoulis. Hence the overall shape is preserved. This drift–fluctuation decomposition mirrors classical treatments of stochastic processes vankampen_stochastic. Equation (31) therefore shows that bias does not deform the distribution: it introduces no broadening, skewness, or change in shape. It generates only a uniform drift proportional to the geodesic distance and the imbalance (2r1)(2r-1), while the statistical structure of the underlying fluctuations remains invariant.
@@ VII.3 Consequences for misinformation
so the only influence of $r$ enters through the translation $\Delta\mu_{ij}$. When the true distribution $P(x)$ is Gaussian, this dependence reduces to a simple quadratic function of the drift, since a translation of a Gaussian changes only its mean. Thus, varying the bias probability cannot generate any new qualitative behaviour in the misinformation. It shifts all perceived distributions coherently while the shapes and dispersions remain unchanged. This decomposition into a deterministic drift and a shape-preserving fluctuation term mirrors classical results for biased random walks and additive binomial noise feller_prob; vankampen_stochastic; cover_2006. Similarly, in our network setting, the bias induces only a uniform shift, without altering the form of the perceived distribution.
@@
A key physical implication is that biased errors do not increase the uncertainty of the information flowing through the network. Instead, they produce a coherent, topology-dependent displacement of the perceived state. Consequently, systematic misinformation accumulates along longer paths, making agents consistently incorrect rather than increasingly uncertain. This mechanism aligns with phenomena observed in social systems, where groups may converge to shared but biased narratives while maintaining internal confidence delvicario2016; sunstein2009. Similar effects appear in technological and biological settings—in distributed sensing, small measurement biases can shift global consensus without increasing estimated noise olfatisaber2007, and in regulatory or signaling networks, biased transmission can lead to coherent misregulation rather than merely noisier responses.
@@ VIII Discussion & Conclusions
In ER networks, misinformation displays a distinctive double-peaked profile as the link probability varies. Near the percolation threshold, long geodesic chains amplify accumulated fluctuations, whereas at slightly higher densities the coexistence of extended paths and emerging cycles produces a second, weaker maximum newman_2010; barrat2008. Only at high densities—where shortest paths collapse logarithmically—does misinformation sharply decrease. This behaviour highlights a general principle: misinformation is maximized when the network is sufficiently connected to propagate signals widely yet sparse enough to preserve long, error-amplifying routes.
@@
WS networks reveal a complementary trend. A small amount of rewiring creates a narrow small-world regime in which misinformation is minimized. Sparse random shortcuts reduce path lengths dramatically while retaining strong local structure watts_nature_1998; strogatz2001. This structural crossover resembles transport optimisation in disordered media dorogovtsev2008. Notably, empirical biological networks—neural, metabolic, genetic, and ecological—often fall precisely within this low-to-intermediate randomness range sporns_neuro_2004; jeong2000; alon2007introduction, aligning with our observation that small-world architectures suppress global distortion.
@@
BA networks exhibit the lowest level of misinformation both in the sparse and intermediate-density connectivity regimes. High-degree hubs act as robust integrators that shorten effective distances, dilute noise, and stabilise information flow barabasi_science_1999; albert2000error; jeong_nature_2001; han2004evidence. This result provides a quantitative explanation for the resilience of scale-free systems—heterogeneity and hub dominance inherently suppress the accumulation of pathwise errors. Although BA networks are known to facilitate rapid propagation of both information and errors due to their hub-dominated architecture barthelemy_prl_2004; wang_rsos_2018, our results show that, within a static error-accumulation model, they exhibit robustness to path induced errors. This apparent contrast between quick dynamical error propagation and suppressed misinformation is a nontrivial feature of scale-free topology. This feature highlights an interesting phenomenon that merits investigation in future work.
@@
A central analytical contribution of this work is the demonstration that directional bias in the error process creates only a uniform drift in the perceived state, without altering the fluctuation structure. The biased process splits into an additive drift term and a shape-preserving noise term, similar to classical stochastic-process theory feller_prob; vankampen_stochastic. Consequently, biased errors produce coherent misinformation rather than uncertainty. Our formalism connotes several observed phenomena in social systems, where groups converge to internally confident but biased narratives sunstein2009; delvicario2016—and in technological or biological networks, where small sensor biases imbalances drive global misregulation olfatisaber2007.
@@
Our framework opens several avenues for further work. Extending the model to weighted, directed, temporal, or multilayer networks would allow the analysis of systems with richer dynamical features. Incorporating correlated or memory-dependent errors could reveal how temporal persistence modifies misinformation accumulation. From an optimisation perspective, one could search for cost-constrained rewiring strategies that will minimize global distortion or identify network modifications that maximize robustness under biased or adversarial noise. Finally, coupling this framework with agent-based dynamics or adaptive topology could shed light on how real systems evolve to balance efficiency, robustness, and misinformation suppression.
Overall, the theoretical and numerical results presented here lay a foundation for a statistical physics of information distortion in complex networks, providing both mechanistic insight and a versatile analytical toolset for future studies.
@@ Acknowledgements
$S$ acknowledges the support of the Department of Chemical and Biological Physics and AMOS at the Weizmann Institute of Science, Israel, where a significant part of this work was carried out. JCC acknowledges financial support from the EPSRC through grant EP/V048740/2. AM acknowledges support from the School of Computer Science, University of Birmingham, and the Centre for Ecological Sciences, Indian Institute of Science, Bengaluru, where parts of this work were conducted. His stay at the University of Birmingham was supported by the EPSRC through grant EP/V048740/2, and his stay at the Indian Institute of Science, Bengaluru, was supported by the SERB (DST, Government of India) through project no. PDF/2023/001151. AM is currently employed at M. S. Ramaiah University of Applied Sciences. RC gratefully acknowledges the generous allocation of computing resources by the Department of Theoretical Physics (DTP), Tata Institute of Fundamental Research (TIFR), along with related technical assistance from Kapil Ghadiali and Ajay Salve.
@@ Appendix A Mean-field estimate of misinformation for a Gaussian true distribution
where $\overline{m^{k}}=\int_{c=0}^{1}(\alpha c)^{k}\rho(c)\,dc$. Defining $\Delta(x)=Q(x)-P(x)$ yields

### v3

@@ VI.3 Barabási–Albert Networks
Scale-free networks generated by the Barabási–Albert (BA) preferential attachment model Barabási and Albert (1999) provide a natural framework for examining the role of heterogeneous connectivity in misinformation. In this model, each newly added node attaches to $m$ existing nodes with probability proportional to their degrees, producing a power-law degree distribution
@@
The biological relevance of scale-free architecture is well documented. Metabolic networks Jeong et al. (2000); Wagner and Fell (2001), gene-regulatory circuits Lee et al. (2002); Babu et al. (2004), and protein–protein interaction networks Jeong et al. (2001); Han et al. (2004) all display degree exponents between 2 and 3. In such systems, hubs frequently correspond to essential components—such as ATP in metabolism or master transcriptional regulators—that coordinate activity across multiple pathways and support robust communication. For context, Table 2 compares several biological scale-free networks with a BA network of matched average degree. Although empirical exponents vary across the range 2–3, the BA model consistently produces an exponent close to 3.
@@ VII.2 Translation of the perceived distribution
To understand why Eq. (31) preserves the functional shape of the distribution, consider the case where the true state $X_{i}$ is Gaussian. The perceived value $X^{(j\leftarrow i)}$ is the sum of this Gaussian variable, a constant drift, and a centered binomial fluctuation. Because the fluctuation term is a sum of independent Bernoulli increments, it is approximately Gaussian by the central limit theorem Feller (1968); and the sum of two Gaussian variables is Gaussian Port (1986). Hence the overall shape is preserved. This drift–fluctuation decomposition mirrors classical treatments of stochastic processes Van Kampen (2008). Equation (31) therefore shows that bias does not deform the distribution: it introduces no broadening, skewness, or change in shape. It generates only a uniform drift proportional to the geodesic distance and the imbalance (2r1)(2r-1), while the statistical structure of the underlying fluctuations remains invariant.
@@ VII.3 Consequences for misinformation
so the only influence of $r$ enters through the translation $\Delta\mu_{ij}$. When the true distribution $P(x)$ is Gaussian, this dependence reduces to a simple quadratic function of the drift, since a translation of a Gaussian changes only its mean. Thus, varying the bias probability cannot generate any new qualitative behaviour in the misinformation. It shifts all perceived distributions coherently while the shapes and dispersions remain unchanged. This decomposition into a deterministic drift and a shape-preserving fluctuation term mirrors classical results for biased random walks and additive binomial noise Feller (1968); Van Kampen (2008); Cover and Thomas (2006). Similarly, in our network setting, the bias induces only a uniform shift, without altering the form of the perceived distribution.
@@
A key physical implication is that biased errors do not increase the uncertainty of the information flowing through the network. Instead, they produce a coherent, topology-dependent displacement of the perceived state. Consequently, systematic misinformation accumulates along longer paths, making agents consistently incorrect rather than increasingly uncertain. This mechanism aligns with phenomena observed in social systems, where groups may converge to shared but biased narratives while maintaining internal confidence Vicario et al. (2016a); Sunstein (2009). Similar effects appear in technological and biological settings—in distributed sensing, small measurement biases can shift global consensus without increasing estimated noise Olfati-Saber et al. (2007), and in regulatory or signaling networks, biased transmission can lead to coherent misregulation rather than merely noisier responses.
@@ VIII Discussion & Conclusions
In ER networks, misinformation displays a distinctive double-peaked profile as the link probability varies. Near the percolation threshold, long geodesic chains amplify accumulated fluctuations, whereas at slightly higher densities the coexistence of extended paths and emerging cycles produces a second, weaker maximum Newman (2010); Barrat et al. (2008). Only at high densities—where shortest paths collapse logarithmically—does misinformation sharply decrease. This behaviour highlights a general principle: misinformation is maximized when the network is sufficiently connected to propagate signals widely yet sparse enough to preserve long, error-amplifying routes.
@@
WS networks reveal a complementary trend. A small amount of rewiring creates a narrow small-world regime in which misinformation is minimized. Sparse random shortcuts reduce path lengths dramatically while retaining strong local structure Watts and Strogatz (1998); Strogatz (2001). This structural crossover resembles transport optimisation in disordered media Dorogovtsev et al. (2008). Notably, empirical biological networks—neural, metabolic, genetic, and ecological—often fall precisely within this low-to-intermediate randomness range Sporns and Zwi (2004); Jeong et al. (2000); Alon (2007), aligning with observation that small-world architectures suppress global distortion.
@@
BA networks exhibit the lowest level of misinformation both in the sparse and intermediate-density connectivity regimes. High-degree hubs act as robust integrators that shorten effective distances, dilute noise, and stabilise information flow Barabási and Albert (1999); Albert et al. (2000); Jeong et al. (2001); Han et al. (2004). This result provides a quantitative explanation for the resilience of scale-free systems—heterogeneity and hub dominance inherently suppress the accumulation of pathwise errors. Although BA networks are known to facilitate rapid propagation of both information and errors due to their hub-dominated architecture Barthélemy et al. (2004); Wang et al. (2018), our results show that, within a static error-accumulation model, they exhibit robustness to path induced errors. This apparent contrast between quick dynamical error propagation and suppressed misinformation is a nontrivial feature of scale-free topology. This feature highlights an interesting phenomenon that merits investigation in future work.
@@
A central analytical contribution of this work is the demonstration that directional bias in the error process creates only a uniform drift in the perceived state, without altering the fluctuation structure. The biased process splits into an additive drift term and a shape-preserving noise term, similar to classical stochastic-process theory Feller (1968); Van Kampen (2008). Consequently, biased errors produce coherent misinformation rather than uncertainty. Our formalism connotes several observed phenomena in social systems, where groups converge to internally confident but biased narratives Sunstein (2009); Vicario et al. (2016a)—and in technological or biological networks, where small sensor biases imbalances drive global misregulation Olfati-Saber et al. (2007).
@@
The framework introduced in our work is deliberately static, or asymptotic in time: information is assumed to propagate along shortest paths in a single pass, without any time dependence, feedback loops, memory effects, or multi-round updating. This abstraction deviates from many real-world misinformation propagation processes, where delays, competition between narratives, repeated exposures, and corrective mechanisms play central roles Centola (2010); Vicario et al. (2016b); Vosoughi et al. (2018); Newman (2018). Our objective, however, is not to reproduce the full temporal complexity of social or biological communication, but rather to quantify the contribution of network topology to error propagation under minimal and well-controlled dynamical assumptions. In this sense, shortest-path propagation provides a natural assumption that captures the most direct routes of error propagation without losing the interpretation of the role of connectivity. More general schemes, such as diffusion-based spreading, averaging over multiple paths, or incorporating adaptive, or feedback-driven dynamics, would enrich the model, but at the cost of increased complexity. We therefore view the present framework as a foundational baseline rather than a comprehensive analysis of the error spreading process. We emphasize that our purpose is to establish an analytically tractable limit on misinformation. Such limits can serve as a reference point for future realistic extensions incorporating temporal evolution, alternative routing mechanisms, and heterogeneous agent behavior.
Our model assumes that each transmission introduces an independent error drawn from a common distribution. We adopt this assumption to ensure analytical tractability and to isolate the influence of network topology on error propagation. Within this assumption, our framework already accommodates certain forms of heterogeneity, including asymmetric binomial noise, and Gaussian initial state distributions, thereby demonstrating that the qualitative behavior of error propagation is robust to changes in the distribution-shape, variance, and skewness. Nevertheless, the present framework does not capture correlated noise, either temporal or spatial, and source-dependent biases, such as misinformation introduced by some influential nodes, coordinated media outlets, or tightly coupled communities Mezard et al. (1986); Lazer et al. (2018); Vicario et al. (2016a). Allowing node-specific error distributions, correlated transmissions, or adversarial sources could give rise to qualitatively new phenomena, including localization of error, amplification by central actors, or nontrivial spatial clustering. Exploring such effects would require relaxing the assumption of independence and extending the framework to incorporate heterogeneous, correlated, or memory-dependent noise processes Gleeson et al. (2016); Pastor-Satorras et al. (2015). We leave these directions for future work, while we emphasize that the present framework establishes a baseline upon which more elaborate models can be built.
Since the KL divergence measures the mismatch between a true distribution and its distorted form, it turns out to be the simplest yet standard choice in our framework to measure the misinformation Kullback and Leibler (1951); Cover and Thomas (2006). In principle, one can quantify misinformation between a perceived distribution and the true distribution using a more general class of asymmetric divergence measures, namely the Rényi divergences. The Rényi divergences form a one-parameter family of divergence measures indexed by an order parameter $\alpha>0$, $\alpha\neq 1$. The KL divergence appears as a special (limiting) case of the Rényi divergence in the limit $\alpha\to 1$ Rényi (1961). More generally, Rényi divergences of different orders emphasize different regions of the distributions: values $\alpha<1$ give more weight to rare events and distribution tails, whereas values $\alpha>1$ place greater weight on high-probability events Cover and Thomas (2006); van Erven and Harremos (2014). However, as shown in App. B, in certain limits the Rényi divergences are related to the KL divergence in a simple manner. Specifically, (i) when the true-state probability density function is a Dirac-delta distribution, the Rényi divergences reduce to the KL divergence, irrespective of the value of $\alpha$, implying the misinformation remains unchanged; (ii) when the true distribution is sufficiently broad, the Rényi divergences become proportional to the KL divergence, with proportionality factor $\alpha$, implying that the misinformation is trivially scaled. Thus in both these cases, our main results including the relative ranking of network topologies and the presence of crossover behavior remain unchanged. For more general cases, the impact of $\alpha$ on the main qualitative trends, such as the relative ranking of network topologies and the presence of crossover behavior, remains an interesting question and is left for future investigation.
@@
Overall, our theoretical and numerical results lay a foundation for a statistical model of information distortion in complex networks, providing both insight and a versatile analytical tool for future studies. In addition to the directions mentioned above, extending the model to weighted, directed, or multilayer networks would allow the analysis of systems with richer dynamical features. Incorporating correlated or memory-dependent errors could reveal how temporal persistence modifies misinformation propagation. From an optimization perspective, one could also search for cost-constrained rewiring strategies that will minimize global distortion or identify networks that maximize the robustness under biased noise. Finally, coupling this framework with agent-based dynamics or adaptive topology could shed light on how real systems evolve to balance efficiency, robustness, and misinformation suppression.
@@ Acknowledgements
SS acknowledges the support of the Department of Chemical and Biological Physics and AMOS at the Weizmann Institute of Science, Israel, where a significant part of this work was carried out. JCC acknowledges financial support from the EPSRC through grant EP/V048740/2. AM acknowledges support from the School of Computer Science, University of Birmingham, and the Centre for Ecological Sciences, Indian Institute of Science, Bengaluru, where parts of this work were conducted. His stay at the University of Birmingham was supported by the EPSRC through grant EP/V048740/2, and his stay at the Indian Institute of Science, Bengaluru, was supported by the SERB (DST, Government of India) through project no. PDF/2023/001151. AM is currently employed at M. S. Ramaiah University of Applied Sciences. RC gratefully acknowledges the generous allocation of computing resources by the Department of Theoretical Physics (DTP), Tata Institute of Fundamental Research (TIFR), along with related technical assistance from Kapil Ghadiali and Ajay Salve.
@@ Appendix A Mean-field estimate of misinformation for a Gaussian true distribution
where $\overline{m^{k}}=\int_{c=0}^{1}(\alpha c)^{k}\rho(c)\,dc$. Therefore, $\Delta(x)=Q(x)-P(x)$ yields
@@
Appendix B Relation between Rényi and KL divergences in special limits
Theorem 1
Let $P(x)=\delta(x=0)$ be the Dirac probability measure peaked at $x=0$, and let $Q(x)$ be a probability measure such that $Q(x=0)>0$, so that $P\ll Q$. Then for all $\alpha>0$, $\alpha\neq 1$, the Rényi divergence satisfies $D_{\alpha}(P\|Q)=D_{\mathrm{KL}}(P\|Q)$ and this equality holds for all $\alpha>0$ under the assumption $P\ll Q$. Note: The notation $P\ll Q$ between two probability distributions signifies that $P$ is absolutely continuous with respect to $Q$.
Proof: We follow the definition and conventions of Rényi divergence given in Erven et al. van Erven and Harremos (2014). Since $P\ll Q$, the Radon–Nikodym derivative $dP/dQ$ exists and is given by Shiryaev (1996)
$\frac{dP(x)}{dQ(x)}=\begin{cases}\dfrac{1}{Q(x=0)},&x=0,\\[6.0pt] 0,&x\neq 0.\end{cases}$ (46)
For all $\alpha>0$, $\alpha\neq 1$, the Rényi divergence is defined by integration with respect to $Q$ as van Erven and Harremos (2014)
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\int\left(\frac{dP}{dQ}\right)^{\alpha}dQ.$ (47)
Substituting the explicit form of $dP/dQ$, we obtain
$\int\left(\frac{dP}{dQ}\right)^{\alpha}dQ=\int_{\{0\}}\left(\frac{1}{Q(x=0)}\right)^{\alpha}dQ=Q(x=0)^{1-\alpha}.$ (48)
Therefore,
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\left(Q(x=0)^{1-\alpha}\right)=-\log Q(x=0).$ (49)
This expression is independent of $\alpha$.
Finally, the KL divergence is given by
$D_{\mathrm{KL}}(P\|Q)$ $=$ $\int\log\left(\frac{dP}{dQ}\right)dP=-\log Q(x=0).$
Hence, $D_{\alpha}(P\|Q)=D_{\mathrm{KL}}(P\|Q)$ for all $\alpha>0$, $\alpha\neq 1$, under the assumption $P\ll Q$.
Theorem 2
Let $P(x)$ and $Q(x)$ be smooth, strictly positive PDFs, then for any $\alpha>0$, $\alpha\neq 1$, the Rényi divergence satisfies $D_{\alpha}(P\|Q)=\alpha\,D_{\mathrm{KL}}(P\|Q)+\mathcal{O}(\Delta^{3})$.
Proof: The Rényi divergence of order $\alpha$ is defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\int dx\,P(x)^{\alpha}Q(x)^{1-\alpha}.$ (51)
Substituting $Q(x)=P(x)+\Delta(x)$ gives
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\int dx\,P(x)\left(1+\frac{\Delta(x)}{P(x)}\right)^{1-\alpha}.$ (52)
Using the expansion
$(1+u)^{1-\alpha}=1+(1-\alpha)u-\frac{\alpha(1-\alpha)}{2}u^{2}+\mathcal{O}(u^{3})$, we obtain
$\int dxP(x)\left(1+\frac{\Delta(x)}{P(x)}\right)^{1-\alpha}$ (53)
$=$ $\int dx\Big(P(x)+(1-\alpha)\Delta(x)-\frac{\alpha(1-\alpha)}{2}\frac{\Delta(x)^{2}}{P(x)}$
$+\mathcal{O}\left(\frac{|\Delta(x)|^{3}}{P(x)^{2}}\right)\Big).$
The linear term vanishes due to normalization, $\int dx~\Delta(x)=0$. Therefore the above expression becomes
$\int dx~P(x)\left(1+\frac{\Delta(x)}{P(x)}\right)^{1-\alpha}$ (54)
$=$ $1-\frac{\alpha(1-\alpha)}{2}\int dx\left[\frac{\Delta(x)^{2}}{P(x)}+\mathcal{O}\left(\frac{|\Delta(x)|^{3}}{P(x)^{2}}\right)\right].$
Taking the logarithm of 54 and expanding the logarithm in a Taylor series, and retaining terms up to $|\Delta|^{2}$, we obtain
$D_{\alpha}(P\|Q)=\frac{\alpha}{2}\int dx\,\frac{\Delta(x)^{2}}{P(x)}+\mathcal{O}\left(\int dx~\frac{|\Delta(x)|^{3}}{P(x)^{2}}\right).$ (55)
On the other hand, from (39)
$D_{\mathrm{KL}}(P\|Q)=\frac{1}{2}\int dx\,\frac{\Delta(x)^{2}}{P(x)}+\mathcal{O}\left(\int dx~\frac{|\Delta(x)|^{3}}{P(x)^{2}}\right).$ (56)
Comparing the two expressions (55) and (56) we show
$D_{\alpha}(P\|Q)=\alpha D_{\mathrm{KL}}(P\|Q)+\mathcal{O}\left(\int dx~\frac{|\Delta(x)|^{3}}{P(x)^{2}}\right).$ (57)